OpenStudy (anonymous):

Find the initial value problem of: dy/dx - 2xy = 2xe^(x^2)

OpenStudy (sidsiddhartha):

are u familiar with linear differential equations of the form \[\frac{ dy }{ dx }+Py=Q\]

OpenStudy (sidsiddhartha):

we can solve if by finding IF(integrating factor)

OpenStudy (anonymous):

I tried to solve it that way but i'm stuck.. Can you show me how to solve this problem?

OpenStudy (sidsiddhartha):

yeah \[\Large \frac{ dy }{ dx }+(-2x)y=2xe^{x^2}\] i can write it in this way okay?

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

so now comapring with the above equation we can write \[P=-2x\]and \[Q=2xe^{x^2}\] right!!

OpenStudy (anonymous):

still correct ;)

OpenStudy (sidsiddhartha):

now are u familiar with this?\[I.F=e^{\int\limits_{}^{}Pdx}\]

OpenStudy (anonymous):

yes i have heard of it but never applied it

OpenStudy (sidsiddhartha):

okay so here we need to apply this method using integrating factor so \[\Large I.F=e^{\int\limits_{}^{}Pdx}=e^{\int\limits_{}^{}-2xdx}=e^{-x^2}\] okay?

OpenStudy (anonymous):

yes! thats what i found

OpenStudy (sidsiddhartha):

al right now consider this equation again\[\frac{ dy }{ dx }+Py=Q\] now after finding I.F ,to solve the diff.equation u have to do this---- \[y*(I.F)=\int\limits_{}^{}Q*(I.F)dx+C\]

OpenStudy (sidsiddhartha):

so we have the \[I.F=e^{-x^2}\] right? now just apply it

OpenStudy (sidsiddhartha):

\[y*(e^{-x^2}=\int\limits_{}^{}2xe^{x^2}*(e^{-x^2})dx+C\] ok getting this?

OpenStudy (anonymous):

yes thats right

OpenStudy (sidsiddhartha):

sorry missed a braket in the lefthand side

OpenStudy (sidsiddhartha):

so now it will be simply \[y*e^{-x^2}=\int\limits_{}^{}2xdx+c\] now i think u can solve it

OpenStudy (sidsiddhartha):

cAn u?

OpenStudy (anonymous):

yes i can! Great. Thank for your help!!!

OpenStudy (sidsiddhartha):

no problem buddy and u can click the "best response" button if u're satisfied with my answer :) have a nice day